OpenStudy (randomecorner):

What is the solution to the equation? 7(5 + 9) = 7 • x + 7 • 9 A. x = 1 B. x = 5 C. x = 7 D. x = 9

OpenStudy (randomecorner):

idk how to do this one

OpenStudy (tkhunny):

This is an eyeball problem. Just look at ti and write down the answer. From the left-hand side 7(5 + 9) = 7*5 + 7*9 From the right-hand side 7 • x + 7 • 9 = 7 • (x + 9) See the parallels?

OpenStudy (randomecorner):

so X is 5?

jagr2713 (jagr2713):

Yep but how did you get it thou ?

OpenStudy (randomecorner):

because he said 7 *5

jagr2713 (jagr2713):

ok

OpenStudy (randomecorner):

know i get it thx that helped ALOT :3thx have a good easter

OpenStudy (randomecorner):

bye